# Kai Shu, Lu Cheng, Huan Liu

### v4

@@ 4.1.2 Tuning Techniques
Supervised fine-tuning (SFT) is the most commonly used approach to facilitate the judge LLMs to learn from pairwise Wang et al. (2024h); Li et al. (2024e); Wang et al. (2023b); Zhu et al. (2023) or pointwise Xie et al. (2024a); Wang et al. (2023b); Yue et al. (2023) judgment data. Among many works that adopted SFT, Vu et al. (2024) propose a supervised multitask training to tune their Foundational Large Autorater Models (FLAMe) across multiple mixed datasets of various tasks. To equip the judge LLM with both pairwise and pointwise judging capabilities, Kim et al. (2024b) novelly propose joint training and weight merging approaches during the tuning stage and find the latter does not improve evaluation performances in the majority of cases. To obtain a judge model that can not only generate responses but also compare pairwise preferences, Lee et al. (2024) devise Judge-augmented Supervised Fine-tuning (JSFT) with an augmented preference learning dataset. During the training phase, Ke et al. (2024) enhance their model by adding simplified prompts to distinguish different parts of inputs and augment pairwise training data by swapping the order of two generated texts and exchanging the corresponding content in critiques.  Xu et al. (2023) further fine-tune their INSTRUCTSCORE model on self-generated outputs to optimize feedback scores, resulting in diagnostic reports that are better aligned with human judgment.  Liu et al. (2024a) also propose a two-stage supervised fine-tuning approach, first applying vanilla instruction tuning to equip the model with the ability to follow instructions for diverse evaluations. They then perform further tuning with auxiliary aspects to enrich the training process, incorporating an additional instruction-tuning stage to leverage potential connections to the target evaluation aspect.
@@
Preference learning is closely aligned with judgment and evaluation tasks, especially comparative and ranking judgment. In addition to works that directly adopt or augment preference learning datasets for supervised fine-tuning judge LLMs, several studies apply preference learning techniques to enhance LLMs’ judging capabilities. To enhance the quality of judgment provided by HALU-J, Wang et al. (2024a) further tune it with Directed Preference Optimization (DPO) Rafailov et al. (2023) after the SFT stage under the multiple-evidence setting. Similarly, Park et al. (2024) apply DPO with synthetic “bad” responses that contain critical errors but exhibit stylistic qualities favored by judge models, helping to mitigate bias in the judge LLMs.  Wu et al. (2024a) novelly propose meta-rewarding, which leverages the policy LLMs to judge the quality of their own judgment and produce pairwise signals for enhancing the LLMs’ judging capability. This concept is also adopted by Wang et al. (2024f), who propose self-taught evaluators that use corrupted instructions to generate suboptimal responses as inferior examples for preference learning. Recently, Hu et al. (2024) propose Themis, an LLM dedicated to NLG evaluation, which has been trained with designed multi-perspective consistency verification and rating-oriented preference alignment methods. Li et al. (2024o) propose PORTIA, an alignment-based approach designed to mimic human comparison behavior to calibrate position bias in an effective manner.
@@ 4.2.2 Rule Augmentation
Rule-augmented prompting involves embedding a set of principles, references, and evaluation rubrics directly within the prompt for the judge LLM. This approach is commonly employed in LLM-based evaluations, where judge LLMs are guided to assess specific aspects Li et al. (2024e); Bai et al. (2023a); Yu et al. (2024c); Qian et al. (2024) and provided with detailed rubrics Gao et al. (2023b); Kim et al. ; Wang et al. (2024e); Murugadoss et al. (2024); Li et al. (2024l, h) to ensure a fair comparison. A distinct approach is seen in Liu et al. (2024b), where the judge LLM is prompted to generate its own scoring criteria through in-context learning on a set of few-shot examples. In alignment with LLM-as-a-judge, Bai et al. (2022) first propose to introduce a list of principles (e.g., helpfulness, harmlessness, honesty) for the judge LLM to compare the two candidates more precisely and directionally. Following them, subsequent works Lee et al. (2023, 2024); Guo et al. (2024); Sun et al. (2024a); Beigi et al. (2024a) enhance this principle-driven prompting by incorporating more detailed explanations for each aspect of the principle or rubric. Besides, both Li and Qiu (2023) and Li et al. (2024c) propose to prompt LLMs to retrieve appropriate demonstrations/ knowledge triples based on the candidates’ helpfulness in solving specific problems. To obtain diverse responses from LLMs, Lahoti et al. (2023) prompt multiple LLMs to judge the diversity of each candidate and select the most diverse one for further polishing. Zhang et al. (2024g) propose RevisEval, which leverages the self-correction capabilities of LLMs to adaptively revise the response, then treat the revised text as the principles for the subsequent evaluation. Recently,some works investigate the reliability of LLM-as-a-personalized-judge Dong et al. (2024); Seraj et al. (2024); Kim and Baek (2024), providing persona as a part of principles for LLMs to make personalized judgments.
@@ 4.2.3 Multi-agent Collaboration
Accessing results from a single LLM judge may not be reliable due to various biases inherent in LLMs Wang et al. (2023c); Liusie et al. (2024); Zhu et al. (2023); Koo et al. (2023a); Liu et al. (2023c). To address this limitation, Li et al. (2023a) introduced the Peer Rank (PR) algorithm, which considers each peer LLM’s pairwise preferences for all answer pairs and produces a final ranking of models. Building on this foundation, several architectures and techniques for multi-agent LLMs have emerged, including mixture-of-agent Zhang et al. (2023b), role play Wu et al. (2023), debating Chan et al. (2023); Zhang et al. (2024f), and voting Zhu et al. (2024c); Rahmani et al. (2024). Jung et al. (2024) proposed Cascaded Selective Evaluation, where less expensive models serve as initial judges, escalating to stronger models only when necessary Beigi et al. (2024a). Additionally, some works apply multi-agent collaboration for alignment data synthesis, leveraging multiple LLM judges to refine responses Arif et al. (2024) or provide more accurate pairwise feedback Li et al. (2024i). Recnetly, Li et al. (2024m) propose the MATEval, where all agents are played by LLMs like GPT-4. The MATEval framework emulates human collaborative discussion methods, integrating multiple agents’ interactions to evaluate open-ended text.
@@ 4.2.5 Multi-turn Interaction
In evaluation, a single response may not provide enough information for an LLM judge to thoroughly and fairly assess each candidate’s performance. To address this limitation, multi-turn interactions are commonly adopted to offer a more comprehensive evaluation. Typically, the process begins with an initial query or topic, followed by dynamically interacting between the judge LLM and candidate models. Bai et al. (2023b) propose a multi-round setting where the evaluator takes on an interviewer role, posing increasingly sophisticated follow-up questions based on the models’ previous answers. Similarly, Yu et al. (2024c) introduce KIEval, a Knowledge-grounded Interactive Evaluation framework, which novelly incorporates an LLM-powered interactor to enable dynamic, contamination-resilient assessments. Additionally, some approaches facilitate debates among candidates in a multi-round format. For example, Zhao et al. (2024c) design a framework where two LLMs engage in a multi-round peer battle around a query, allowing their true performance differences to surface. Moniri et al. (2024) propose an automated benchmarking system where LLMs debate, with the final assessment carried out by another LLM judge.
@@ 5.1.1 Open-ended Generation Tasks
Open-ended generation refers to tasks where the generated content is expected to be safe, accurate, and contextually relevant, though there isn’t a single “correct” answer. Such tasks include dialog response generation, summarization, story generation, and creative writing Badshah and Sajjad (2024); Bai et al. (2023a); Kumar et al. (2024); Zeng et al. ; Yang et al. (2024). Unlike conventional metrics-based evaluation methods, LLM-as-a-judge provides a more nuanced, adaptable, and customized evaluation. As noted by Zheng et al. (2023), LLMs like GPT-4 perform comparably to humans when judging open-ended text generation. In practice, LLM-as-a-judge has been applied to evaluate outputs from a single model to compare outputs from multiple models in a competitive setting. For example, Gao et al. (2023b) use ChatGPT to perform human-like summarization evaluation. Similarly, Wu et al. (2023) propose a comparison-based framework to let LLMs act as judges with multiple role-playing to evaluate the summarization quality in a specific dimension and generate its evaluations.
@@ 5.1.2 Reasoning Tasks
The reasoning capability of LLMs can be evaluated through their intermediate thinking process and final answers on specific reasoning tasks Mondorf and Plank (2024). Recently, LLM-as-a-judge has been used to evaluate the logical progression, depth, and coherence of the model’s intermediate reasoning paths. For mathematical reasoning tasks, Xia et al. (2024) introduce an automatic evaluation framework using a judge LLM specifically designed to assess the quality of reasoning steps in problem-solving processes. LLM-as-a-judge can also be applied to more complex reasoning tasks like temporal reasoning, where models need to understand the relationships of different events over time. Fatemi et al. (2024) build synthetic datasets specifically tailored to evaluate LLMs’ temporal reasoning abilities across varied scenarios, testing their proficiency in reasoning with sequences, causality, and dependencies among temporally ordered events.
@@ 5.1.3 Emerging Tasks
Another line of research has been directed towards evaluating Large Multimodal Models (LMMs) and Large Vision-Language Models (LVLMs). For example, Xiong et al. (2024b) explore LMM-as-a-judge to evaluate the performance of multimodal models, providing both a final score and the underlying rationale for evaluations, promoting transparency and consistency. Chen et al. (2024c) propose the first benchmark for the automatic evaluation of LVLMs, specifically for self-driving corner cases. They find that evaluations performed by LLMs-as-judges align more closely with human preferences than those conducted by LVLM-as-judges.
Recently, we have seen more customized utilization of LLM-as-a-judge to evaluate emerging tasks such as code understanding Zhao et al. (2024a), legal knowledge Fei et al. (2023), game development Isaza-Giraldo et al. (2024), ocean science Bi et al. (2023), healthcare conversations Wang et al. (2024j), debating judgment Liang et al. (2024a), retrival-augmented generation Dhole et al. (2024), cross ability Zhong et al. (2024), HCI Li et al. (2024j), role-playing Zhou et al. (2024b), RAG Jin et al. (2024) and more. This trend reflects the growing adaptability of LLMs-as-judges in evaluating diverse and specialized domains.
@@ 5.2.1 Larger Models as Judges
One intuitive idea for adopting LLM-as-a-judge in alignment tuning is to leverage the feedback from larger, more powerful LLMs to guide smaller, less capable models. Bai et al. (2022) first propose employing the AI’s feedback to build a harmless AI assistant. They train the reward model using synthetic preference data based on the preference of a pre-trained language model. Building on this, Lee et al. (2023) discover that the RLAIF method can achieve comparable performance with RLHF even when the LLM judge is not strong enough. They also introduced DIRECT-RLAIF, which directly employs an off-the-shelf LLM as the judge model to mitigate reward staleness in reward models. To avoid reward hacking in alignment, Sun et al. (2024a) devise an instructable reward model trained on synthetic preference data. It enables humans to perform RL-time interventions to better align the target policy with human values. Apart from the abovementioned studies, Guo et al. (2024) introduce online AI feedback (OAIF), directly utilizing the preference signals from an annotation model to train the target model. There are also works that utilize multi-agent cooperation for better judgment in alignment tuning. Arif et al. (2024) and Sengupta et al. (2024) construct a synthetic preference optimization dataset using multi-agent workflows and adopt LLMs as judges with diverse prompting strategies and pipelines. Similarly, Li et al. (2024i) employ multiple LLMs to debate with each other, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To align the generated code with human preference, Weyssow et al. (2024) introduce CodeUltraFeedback, a preference coding dataset constructed using the LLM-as-a-Judge methodology. This synthetic dataset is later used to fine-tune and align small code LLMs using SFT and DPO. Recently, Wang et al. (2024d) propose BPO, using GPT-4 as the judge and constructing synthetic pairwise feedback for knowledge depth and breadth balance in the alignment process.
@@ 5.2.2 Self-Judging
Another line of work aims to utilize the preference signal from the same LLM for self-improving. Yuan et al. (2024c) first propose the concept of a self-rewarding LLM, where pairwise data is constructed by having the LLM itself serve as the judge. Following them, Wu et al. (2024a) introduce meta-rewarding that judges the LLMs’ judge and uses the feedback to refine their judging skills. Their LLM-as-a-meta-judge approach significantly enhances the model’s capacity to evaluate and follow instructions. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. Lee et al. (2024) devise Judge augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. To fully utilize this judge model, they also propose the Self-Rejection by Tournament method to select the best response in the inference time. Unlike the above approaches, which use the LLM as a judge for pairwise data construction, Tong et al. (2024) apply LLM-as-a-judge as a self-filtering method to ensure the quality of synthetic data pairs in alignment tasks for reasoning. To reduce the computation overhead in pairwise judgment, Zhai et al. (2024) propose a ranked pairing method for self-preferring language models, which accelerates the comparison process by measuring the strength of each response against a baseline. Liu et al. (2024c) introduce meta-ranking, enabling weaker LLMs Sun et al. (2024b) to serve as reliable judges and provide trustworthy feedback. They also apply their meta-ranking method in post-SFT training, combining it with Kahneman-Tversky Optimization (KTO) for improved alignment. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024c) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration. Recently, Yasunaga et al. (2024) propose to combine LLM-as-a-judge and data synthesis and successfully build human-aligned LLMs with a few annotation.

### v5

@@ 4.1.2 Tuning Techniques
Supervised fine-tuning (SFT) is the most commonly used approach to facilitate the judge LLMs to learn from pairwise Wang et al. (2024j); Li et al. (2024e); Wang et al. (2023b); Zhu et al. (2023) or pointwise Xie et al. (2024a); Wang et al. (2023b); Yue et al. (2023) judgment data. Among many works that adopted SFT, Vu et al. (2024) propose a supervised multitask training to tune their Foundational Large Autorater Models (FLAMe) across multiple mixed datasets of various tasks. To equip the judge LLM with both pairwise and pointwise judging capabilities, Kim et al. (2024b) novelly propose joint training and weight merging approaches during the tuning stage and find the latter does not improve evaluation performances in the majority of cases. To obtain a judge model that can not only generate responses but also compare pairwise preferences, Lee et al. (2024) devise Judge-augmented Supervised Fine-tuning (JSFT) with an augmented preference learning dataset. During the training phase, Ke et al. (2024) enhance their model by adding simplified prompts to distinguish different parts of inputs and augment pairwise training data by swapping the order of two generated texts and exchanging the corresponding content in critiques.  Xu et al. (2023) further fine-tune their INSTRUCTSCORE model on self-generated outputs to optimize feedback scores, resulting in diagnostic reports that are better aligned with human judgment.  Liu et al. (2024a) also propose a two-stage supervised fine-tuning approach, first applying vanilla instruction tuning to equip the model with the ability to follow instructions for diverse evaluations. They then perform further tuning with auxiliary aspects to enrich the training process, incorporating an additional instruction-tuning stage to leverage potential connections to the target evaluation aspect.
@@
Preference learning is closely aligned with judgment and evaluation tasks, especially comparative and ranking judgment. In addition to works that directly adopt or augment preference learning datasets for supervised fine-tuning judge LLMs, several studies apply preference learning techniques to enhance LLMs’ judging capabilities. To enhance the quality of judgment provided by HALU-J, Wang et al. (2024a) further tune it with Directed Preference Optimization (DPO) Rafailov et al. (2023) after the SFT stage under the multiple-evidence setting. Similarly, Park et al. (2024a) apply DPO with synthetic “bad” responses that contain critical errors but exhibit stylistic qualities favored by judge models, helping to mitigate bias in the judge LLMs.  Wu et al. (2024a) novelly propose meta-rewarding, which leverages the policy LLMs to judge the quality of their own judgment and produce pairwise signals for enhancing the LLMs’ judging capability. This concept is also adopted by Wang et al. (2024h), who propose self-taught evaluators that use corrupted instructions to generate suboptimal responses as inferior examples for preference learning. Recently, Hu et al. (2024) propose Themis, an LLM dedicated to NLG evaluation, which has been trained with designed multi-perspective consistency verification and rating-oriented preference alignment methods. Li et al. (2024o) propose PORTIA, an alignment-based approach designed to mimic human comparison behavior to calibrate position bias in an effective manner.
@@ 4.2.2 Rule Augmentation
Rule-augmented prompting involves embedding a set of principles, references, and evaluation rubrics directly within the prompt for the judge LLM. This approach is commonly employed in LLM-based evaluations, where judge LLMs are guided to assess specific aspects Li et al. (2024e); Bai et al. (2023a); Yu et al. (2024d); Qian et al. (2024) and provided with detailed rubrics Gao et al. (2023b); Kim et al. ; Wang et al. (2024f); Murugadoss et al. (2024); Li et al. (2024l, h); Chen et al. (2024c) to ensure a fair comparison. A distinct approach is seen in Liu et al. (2024b), where the judge LLM is prompted to generate its own scoring criteria through in-context learning on a set of few-shot examples. In alignment with LLM-as-a-judge, Bai et al. (2022) first propose to introduce a list of principles (e.g., helpfulness, harmlessness, honesty) for the judge LLM to compare the two candidates more precisely and directionally. Following them, subsequent works Lee et al. (2023, 2024); Guo et al. (2024); Sun et al. (2024a); Beigi et al. (2024a) enhance this principle-driven prompting by incorporating more detailed explanations for each aspect of the principle or rubric. Besides, both Li and Qiu (2023) and Li et al. (2024c) propose to prompt LLMs to retrieve appropriate demonstrations/ knowledge triples based on the candidates’ helpfulness in solving specific problems. To obtain diverse responses from LLMs, Lahoti et al. (2023) prompt multiple LLMs to judge the diversity of each candidate and select the most diverse one for further polishing. Zhang et al. (2024g) propose RevisEval, which leverages the self-correction capabilities of LLMs to adaptively revise the response, then treat the revised text as the principles for the subsequent evaluation. Recently,some works investigate the reliability of LLM-as-a-personalized-judge Dong et al. (2024); Seraj et al. (2024); Kim and Baek (2024), providing persona as a part of principles for LLMs to make personalized judgments.
@@ 4.2.3 Multi-agent Collaboration
Accessing results from a single LLM judge may not be reliable due to various biases inherent in LLMs Wang et al. (2023c); Liusie et al. (2024); Zhu et al. (2023); Koo et al. (2023a); Liu et al. (2023c). To address this limitation, Li et al. (2023a) introduced the Peer Rank (PR) algorithm, which considers each peer LLM’s pairwise preferences for all answer pairs and produces a final ranking of models. Building on this foundation, several architectures and techniques for multi-agent LLMs have emerged, including mixture-of-agent Zhang et al. (2023b), role play Wu et al. (2023), debating Chan et al. (2023); Zhang et al. (2024f); Feng et al. (2024b), and voting Zhu et al. (2024c); Rahmani et al. (2024). Jung et al. (2024) proposed Cascaded Selective Evaluation, where less expensive models serve as initial judges, escalating to stronger models only when necessary Beigi et al. (2024a). Additionally, some works apply multi-agent collaboration for alignment data synthesis, leveraging multiple LLM judges to refine responses Arif et al. (2024) or provide more accurate pairwise feedback Li et al. (2024i). Recnetly, Li et al. (2024m) propose the MATEval, where all agents are played by LLMs like GPT-4. The MATEval framework emulates human collaborative discussion methods, integrating multiple agents’ interactions to evaluate open-ended text.
@@ 4.2.5 Multi-turn Interaction
In evaluation, a single response may not provide enough information for an LLM judge to thoroughly and fairly assess each candidate’s performance. To address this limitation, multi-turn interactions are commonly adopted to offer a more comprehensive evaluation. Typically, the process begins with an initial query or topic, followed by dynamically interacting between the judge LLM and candidate models. Bai et al. (2023b) propose a multi-round setting where the evaluator takes on an interviewer role, posing increasingly sophisticated follow-up questions based on the models’ previous answers. Similarly, Yu et al. (2024d) introduce KIEval, a Knowledge-grounded Interactive Evaluation framework, which novelly incorporates an LLM-powered interactor to enable dynamic, contamination-resilient assessments. Additionally, some approaches facilitate debates among candidates in a multi-round format. For example, Zhao et al. (2024c) design a framework where two LLMs engage in a multi-round peer battle around a query, allowing their true performance differences to surface. Moniri et al. (2024) propose an automated benchmarking system where LLMs debate, with the final assessment carried out by another LLM judge.
@@ 5.1.1 Open-ended Generation Tasks
Open-ended generation refers to tasks where the generated content is expected to be safe, accurate, and contextually relevant, though there isn’t a single “correct” answer. Such tasks include dialog response generation, summarization, story generation, and creative writing Badshah and Sajjad (2024); Bai et al. (2023a); Kumar et al. (2024); Zeng et al. ; Yang et al. (2024); Bermejo (2024). Unlike conventional metrics-based evaluation methods, LLM-as-a-judge provides a more nuanced, adaptable, and customized evaluation. As noted by Zheng et al. (2023), LLMs like GPT-4 perform comparably to humans when judging open-ended text generation. In practice, LLM-as-a-judge has been applied to evaluate outputs from a single model to compare outputs from multiple models in a competitive setting. For example, Gao et al. (2023b) use ChatGPT to perform human-like summarization evaluation. Similarly, Wu et al. (2023) propose a comparison-based framework to let LLMs act as judges with multiple role-playing to evaluate the summarization quality in a specific dimension and generate its evaluations.
@@ 5.1.2 Reasoning Tasks
The reasoning capability of LLMs can be evaluated through their intermediate thinking process and final answers on specific reasoning tasks Mondorf and Plank (2024); Bernard et al. (2024). Recently, LLM-as-a-judge has been used to evaluate the logical progression, depth, and coherence of the model’s intermediate reasoning paths. For mathematical reasoning tasks, Xia et al. (2024) introduce an automatic evaluation framework using a judge LLM specifically designed to assess the quality of reasoning steps in problem-solving processes. LLM-as-a-judge can also be applied to more complex reasoning tasks like temporal reasoning, where models need to understand the relationships of different events over time. Fatemi et al. (2024) build synthetic datasets specifically tailored to evaluate LLMs’ temporal reasoning abilities across varied scenarios, testing their proficiency in reasoning with sequences, causality, and dependencies among temporally ordered events.
@@ 5.1.3 Emerging Tasks
Another line of research has been directed towards evaluating Large Multimodal Models (LMMs) and Large Vision-Language Models (LVLMs). For example, Xiong et al. (2024b) explore LMM-as-a-judge to evaluate the performance of multimodal models, providing both a final score and the underlying rationale for evaluations, promoting transparency and consistency. Chen et al. (2024d) propose the first benchmark for the automatic evaluation of LVLMs, specifically for self-driving corner cases. They find that evaluations performed by LLMs-as-judges align more closely with human preferences than those conducted by LVLM-as-judges.
Recently, we have seen more customized utilization of LLM-as-a-judge to evaluate emerging tasks such as code understanding Zhao et al. (2024a); Quan et al. (2025); Yu et al. (2024c); Feng et al. (2024a); Lu et al. (2024a); Biagiola et al. (2024), legal knowledge Fei et al. (2023), game development Isaza-Giraldo et al. (2024), ocean science Bi et al. (2023), healthcare conversations Wang et al. (2024m), debating judgment Liang et al. (2024a), retrival-augmented generation Dhole et al. (2024), cross ability Zhong et al. (2024), HCI Li et al. (2024j); Chai et al. (2025); Goren et al. (2024), role-playing Zhou et al. (2024b), RAG Jin et al. (2024), speech synthesis Park et al. (2024b), counterspeech generation Bennie et al. (2025b, a) and more. This trend reflects the growing adaptability of LLMs-as-judges in evaluating diverse and specialized domains.
@@ 5.2.1 Larger Models as Judges
One intuitive idea for adopting LLM-as-a-judge in alignment tuning is to leverage the feedback from larger, more powerful LLMs to guide smaller, less capable models. Bai et al. (2022) first propose employing the AI’s feedback to build a harmless AI assistant. They train the reward model using synthetic preference data based on the preference of a pre-trained language model. Building on this, Lee et al. (2023) discover that the RLAIF method can achieve comparable performance with RLHF even when the LLM judge is not strong enough. They also introduced DIRECT-RLAIF, which directly employs an off-the-shelf LLM as the judge model to mitigate reward staleness in reward models. To avoid reward hacking in alignment, Sun et al. (2024a) devise an instructable reward model trained on synthetic preference data. It enables humans to perform RL-time interventions to better align the target policy with human values. Apart from the abovementioned studies, Guo et al. (2024) introduce online AI feedback (OAIF), directly utilizing the preference signals from an annotation model to train the target model. There are also works that utilize multi-agent cooperation for better judgment in alignment tuning. Arif et al. (2024) and Sengupta et al. (2024) construct a synthetic preference optimization dataset using multi-agent workflows and adopt LLMs as judges with diverse prompting strategies and pipelines. Similarly, Li et al. (2024i) employ multiple LLMs to debate with each other, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To align the generated code with human preference, Weyssow et al. (2024) introduce CodeUltraFeedback, a preference coding dataset constructed using the LLM-as-a-Judge methodology. This synthetic dataset is later used to fine-tune and align small code LLMs using SFT and DPO. Recently, Wang et al. (2024e) propose BPO, using GPT-4 as the judge and constructing synthetic pairwise feedback for knowledge depth and breadth balance in the alignment process.
@@ 5.2.2 Self-Judging
Another line of work aims to utilize the preference signal from the same LLM for self-improving. Yuan et al. (2024e) first propose the concept of a self-rewarding LLM, where pairwise data is constructed by having the LLM itself serve as the judge. Following them, Wu et al. (2024a) introduce meta-rewarding that judges the LLMs’ judge and uses the feedback to refine their judging skills. Their LLM-as-a-meta-judge approach significantly enhances the model’s capacity to evaluate and follow instructions. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. Lee et al. (2024) devise Judge augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. To fully utilize this judge model, they also propose the Self-Rejection by Tournament method to select the best response in the inference time. Unlike the above approaches, which use the LLM as a judge for pairwise data construction, Tong et al. (2024) apply LLM-as-a-judge as a self-filtering method to ensure the quality of synthetic data pairs in alignment tasks for reasoning. To reduce the computation overhead in pairwise judgment, Zhai et al. (2024) propose a ranked pairing method for self-preferring language models, which accelerates the comparison process by measuring the strength of each response against a baseline. Liu et al. (2024c) introduce meta-ranking, enabling weaker LLMs Sun et al. (2024b) to serve as reliable judges and provide trustworthy feedback. They also apply their meta-ranking method in post-SFT training, combining it with Kahneman-Tversky Optimization (KTO) for improved alignment. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024c) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration. Recently, Yasunaga et al. (2024) propose to combine LLM-as-a-judge and data synthesis and successfully build human-aligned LLMs with a few annotation.